# WaveGuard: Robust Deepfake Detection and Source Tracing via Dual-Tree Complex Wavelet and Graph Neural Networks

### v1

@@ III-B Model Architecture
In order to realize a robust watermark embedding and detection system for deepfake, this paper designs a framework, which mainly consists of the following five parts: (1) Frequency Subband Selection Module (FSSM): The frequency subband selection module converts the input image from the spatial domain to the frequency domain and extracts specific frequency sub-band information. (2) Encoder (En): The encoder uses a deep learning network to fuse the specific sub-band information output by the frequency subband selection module with the watermark information, generating the embedded image. (3) Noise Pool Layer (NPL): Simulates common distortions and malicious distortions in actual scenarios to generate the distorted image. (4) Separable Decoder: Includes DensTracer and DensDetector, which are used to extract watermarks under distorted conditions to realize image traceability and detect deepfake, respectively. (5) Structural Consistency Graph Neural Network (SC-GNN): By comparing the image characteristics before and after watermark embedding, the watermark embedding process is optimized to ensure the visual consistency of the watermark image and improve the invisibility of the watermark.
Fig. 4: Overall Architecture of WaveGuard. Frequency Subband Selection Module ($FSSM$) extracts and processes high-frequency sub-bands in the frequency domain, providing essential features for watermark embedding, tracing, and detection. The encoder ($En$) embeds the watermark into the high-frequency sub-bands and reconstructs the watermarked image $I_{w}$. The Noise Pool Layer ($NPL$) applies various distortions to $I_{w}$ to generate distorted images $I_{n}$. The tracer decoder ($Tr$) extracts the watermark $w_{t}$ from $I_{n}$ to evaluate its robustness, while the detector decoder ($De$) extracts the correct watermark $w_{d}$ under common distortions and outputs random-like messages under malicious distortions. In addition, Structural Consistency Graph Neural Network ($SC-GNN$) extracts the structural features of the original image $I_{o}$ and the watermarked image $I_{w}$ by constructing graph representation to ensure the consistency between them.
@@ III-B1 Frequency Subband Selection Module (FSSM)
III-B2 Encoder (En)
@@
The high-frequency sub-band feature output extracted by the frequency subband selection modul undergo Conv-BN-ReLU processing in the Spatial Feature Extraction Module, generating a refined feature map that captures both spatial and fine-grained details. The watermark is transformed into a two-dimensional matrix via a linear layer, resized to match the feature map dimensions, and processed by a convolutional block to extract watermark-specific features. These features are then fused with the shallow spatial features. A three-stage dense connection structure is used for hierarchical watermark embedding. At each stage, the feature map and watermark are concatenated and processed through the DenseBlock Module, which consists of a 1×1 convolution for dimensionality reduction and a 3×3 convolution for feature enhancement. In this process, two different attention mechanisms are adopted: Tanh-based Spatial Embedding Attention Module (TSEAM) and Spatial-wise Image Attention Module (SIMAM). TSEAM is improved on the basis of Convolutional Block Attention Module, which acts on the whole network structure to highlight the important areas in the image. By replacing the ReLU activation function with tanh, the continuity and integrity of watermark features in the [0,1] interval are better preserved. SIMAM is applied after each DenseBlock to weight the local feature map and help the model capture the image details, thus improving the invisibility of the watermark and the robustness to interference. The combination of TSEAM and SIMAM optimizes the watermark embedding process from two dimensions of global features and local details, which not only ensures the effective use of global semantic information, but also realizes the precise control of local features, thus improving the overall visual quality and embedding stability. After the three-stage fusion, the encoder compresses the features using a Down Block, reducing their spatial dimensions while preserving critical information. The final compressed features generate the watermark difference map (wm_diff), which is added to the input image to complete the watermark embedding process. watermark embedding is achieved using the following formula:
$\boldsymbol{U_{\mathit{embed}}=U+\alpha\cdot\mathit{attention\_weight}\cdot\mathit{wm}_{\mathit{diff}}}$ (1)
@@
In Formula (1), $U$ represents the original high-frequency sub-band, $\alpha$ controls the watermark embedding strength, attention_weight is the mask generated by the attention mechanism to weight the watermark embedding area, and $\text{wm}_{\text{diff}}$ is the generated watermark difference map. Through this process, the watermark information is embedded into the original high-frequency sub-band, forming a watermarked high-frequency sub-band $U_{\text{embed}}$. The post-processing module uses the inverse dual-tree complex wavelet transform (IDTCWT) to reconstruct the watermarked U-channel, merges it with the Y and V channels to form watermarked YUV images, and converts them to RGB images $I_{w}$, completing the embedding process.
III-B3 Noise Pool Layer (NPL)
@@ III-B4 Separable Decoder
III-B5 Structural Consistency Graph Neural Network (SC-GNN))
@@ IV-A Implementation Details
Datasets. This research is primarily based on the high-quality CelebA-HQ[39, 40]dataset, which contains 30,000 facial images and 6,217 unique identities. According to the official partitioning, it is divided into training, validation, and testing sets, where the training set contains 24,183 images, and the validation and test sets contain 2,993 and 2,824 images, respectively. To verify the adaptability and robustness of the algorithm across different resolutions and simulate the diverse requirements of practical application scenarios, the experiment uses two types of images with resolutions of 128×128 and 256×256 for testing.
Manipulation Pools. In order to comprehensively evaluate the robustness and adaptability of the algorithm under different distortion conditions, this study constructs two types of operation pools: the Common Pool and Malicious Pool. The Common Pool includes a variety of common image distortion operations, such as {Identity, JpegTest, Resize, GaussianBlur, MedianBlur, Dropout, SaltPepper, GaussianNoise} . These operations are used in the training and testing stages to simulate the general distortions that may occur in real-world scenarios, while ensuring that the facial content of the image remains unchanged. The Malicious Pool includes two common types of deepfake technologies: face swapping and face reenactment. Specifically, SimSwap[2] is used for face swapping, while GANimation[4] and StarGAN[3] are used for face reenactment. In the SimSwap task, the target face is randomly selected from the CelebA verification set. ”In the tasks of GANimation and StarGAN, the target expressions and attributes are randomly selected. The attributes are chosen from the set Male, Young, BlackHair, BlondHair, BrownHair. All forged images are generated using the publicly released pre-trained models to ensure the high credibility and diversity of deepfake images.
Table I
OBJECTIVE VISUAL QUALITY EVALUATION OF WATERMARKED IMAGES ON CELEBA-HQ: MODEL, IMAGE SIZE, MESSAGE LENGTH, PSNR, AND SSIM
@@
Parameters. This research is based on PyTorch[41] and runs on an NVIDIA RTX 4090 GPU. The Adam optimizer[42]is initially set with a learning rate of 0.0005, which is later dynamically adjusted during training. The total loss function is composed of multiple loss terms, which are combined by weighting to balance the influence of each loss on model optimization. As shown in Formula (7), during the training process, the weight parameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to 0.01, 1, 10, 10, and 10, respectively. These weights are adjusted reasonably to reflect the relative importance of each loss term and ensure the convergence and stability of the model in the optimization process.
Table II
COMPARISON OF BIT ERROR RATE IN WATERMARK RECOVERY FOR CELEBA-HQ ACROSS COMMON AND DEEPFAKE DISTORTIONS AT 128$\times$128 AND 256$\times$256 RESOLUTIONS
@@
Baselines. In order to verify the performance of this research method, we selected several representative watermarking and proactive watermarks frameworks as baselines. Although this research belongs to the field of proactive watermarks, due to the limited existing methods and the scarcity of open source codes, we also refer to the robust watermarking method to evaluate its stability and extraction performance. In the watermarking domain, three robust watermarking methods were selected: MBRS [27], CIN [28], and PIMOG [35] (suitable for 128×128 images). In the proactive watermarks domain, we include FaceSigns[16] (for 256×256 images) and SepMark[18] (which supports both 128×128 and 256×256 resolutions). All the other baseline methods use publicly available pre-trained model weights to ensure fairness.
@@ IV-B Visual Quality
To evaluate the visual quality of the image after watermark embedding, this paper adopts two indicators: average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM)[43], which respectively quantify the changes in image signal distortion and structural fidelity, fully reflecting the impact of embedding on the image. Table 1 quantitatively verifies that our method has the best visual quality in all baselines and shows better invisibility. Experimental results show that our method achieves the highest PSNR and SSIM values at resolutions of 128×128 and 256×256. This improvement is primarily attributed to SC-GNN and attention-enhanced embedding. In addition, Figure 5 verifies that there is almost no difference between the watermark image and the original image from a subjective point of view. Combining the quantitative evaluation in Table 1 with the visual comparison in Fig. 5, we can see that our watermark has superior invisibility in both objective indicators and subjective perception.
@@ IV-C Robustness Test
Table 2 quantitatively compares the BER (Bit Error Rate) performance of different methods under common distortion and deepfake forgery, covering 128×128 and 256×256 resolutions. Experimental results show that our method achieves the lowest BER in all test scenarios, showing excellent robustness. Under ordinary distortion, the BER of Tracer and Detector are both close to 0%, which ensures that the watermark can be recovered with high precision under all kinds of interference. However, the BER of the contrast method is obviously higher under various common distortion conditions, which indicates that its robustness is weak. Under the deepfake forgery, the BER of Tracer is still close to 0%, while the BER of semi-robust Detector is maintained close to 50%, which effectively distinguishes the true and forged contents. However, the existing methods have large fluctuations in BER and lack of detection reliability. This robustness is mainly due to the frequency domain embedding strategy. Compared with the spatial domain method, the frequency domain watermark is more stable under common distortion, ensuring the integrity of the watermark information. At the same time, the design of joint sub-band enhances the discrimination ability of semi-robust Detector under deepfake forgery, so that it can stably distinguish real and forged content.
Table III
ABLATION RESULTS FOR PSNR, SSIM, AND TRACER’S BIT ERROR RATE (%), WHERE ‘×’ DENOTES NON-USAGE AND ‘$\sqrt{}$’ DENOTES USAGE.
@@
Table IV
ABLATION RESULTS FOR PSNR, SSIM, AND DETECTOR’S BIT ERROR RATE (%), WHERE ‘×’ DENOTES NON-USAGE AND ‘$\sqrt{}$’ DENOTES USAGE.
@@ IV-D Ablation Study
In order to evaluate the contribution of Joint Sub-bands, Attention Module, Frequency Domain Embedding, and GNN in this method, we conducted ablation experiments on both Tracer and Detector. The experiments were conducted on images with a resolution of 128×128. (a) Adjust the sub-bands usage of Tracer and Detector: In Tracer, watermark embedding was reduced from 1,3,4,6 sub-bands to 1,3, while in Detector, the detection was expanded from 1,3 sub-bands to 1,3,4,6. This allows us to evaluate the influence of joint sub-band information redundancy on watermark recovery and deepfake detection. (b) Remove the attention mechanism in the process of watermark embedding to analyze its influence on visual quality; (c) When the DT-CWT frequency domain embedding is removed, the watermark is directly embedded in the U channel. (d) Removing SC-GNN to evaluate its impact on the visual quality of the watermarked image. Tables 3 and 4 present the results of the ablation study.

### v3

@@ III-B Model Architecture
In order to realize a robust watermark embedding and detection system for deepfake, this paper designs a framework, which mainly consists of the following six parts: (1) Frequency Subband Selection Module (FSSM): The frequency subband selection module converts the input image from the spatial domain to the frequency domain and extracts specific frequency sub-band information. (2) Attention Mechanisms: To enhance watermark imperceptibility, attention mechanisms are introduced to focus on regions more suitable for embedding. By highlighting visually subtle yet semantically important areas, the framework improves watermark invisibility while preserving robustness against distortions. (3) Encoder (En): The encoder uses a deep learning network to fuse the specific sub-band information output by the frequency subband selection module with the watermark information, generating the embedded image. (4) Noise Pool Layer (NPL): Simulates common distortions and malicious distortions in actual scenarios to generate the distorted image. (5) Separable Decoder: Includes DensTracer and DensDetector, which are used to extract watermarks under distorted conditions to realize image traceability and detect deepfake, respectively. (6) Structural Consistency Graph Neural Network (SC-GNN): By comparing the image characteristics before and after watermark embedding, the watermark embedding process is optimized to ensure the visual consistency of the watermark image and improve the invisibility of the watermark.
Fig. 4: Overall Architecture of WaveGuard. Frequency Subband Selection Module ($FSSM$) extracts and processes high-frequency sub-bands in the frequency domain, providing essential features for watermark embedding, tracing, and detection. Attention Module ($AM$) highlights salient regions in the image and guides the features to focus on key areas, thereby enhancing the robustness and effectiveness of watermark embedding. The encoder ($En$) embeds the watermark into the high-frequency sub-bands and reconstructs the watermarked image $I_{w}$. The Noise Pool Layer ($NPL$) applies various distortions to $I_{w}$ to generate distorted images $I_{n}$. The tracer decoder ($Tr$) extracts the watermark $w_{t}$ from $I_{n}$ to evaluate its robustness, while the decoder ($De$) extracts the correct watermark $w_{d}$ under common distortions and outputs random-like messages under malicious distortions. In addition, Structural Consistency Graph Neural Network ($SC-GNN$) extracts the structural features of the original image $I_{o}$ and the watermarked image $I_{w}$ by constructing graph representation to ensure the consistency between them.
@@ III-B1 Frequency Subband Selection Module (FSSM)
III-B2 Attention Mechanisms (AM)
To enhance the imperceptibility of watermark embedding, we introduce two complementary attention mechanisms into the framework. The Tanh-based Spatial Embedding Attention Module (TSEAM) is an improved version of the classic Convolutional Block Attention Module[41], where the ReLU activation function is replaced with Tanh to better accommodate continuous feature representations within the [0,1] range, thereby preserving the continuity and integrity of the watermark. TSEAM operates across the entire network to highlight semantically important yet visually inconspicuous regions, guiding the global embedding strategy. The Spatial-wise Image Attention Module (SIMAM)[42] is applied after each DenseBlock to assign spatial weights to local feature maps, enhancing the model’s sensitivity to fine-grained regions and further improving the stability and robustness of watermark embedding.
III-B3 Encoder (En)
@@
The high-frequency sub-band feature output extracted by the frequency subband selection modul undergo Conv-BN-ReLU processing in the Spatial Feature Extraction Module, generating a refined feature map that captures both spatial and fine-grained details. The watermark is transformed into a two-dimensional matrix via a linear layer, resized to match the feature map dimensions, and processed by a convolutional block to extract watermark-specific features. These features are then fused with the shallow spatial features. A three-stage dense connection structure is used for hierarchical watermark embedding. At each stage, the feature map and watermark are concatenated and processed through the DenseBlock Module, which consists of a 1×1 convolution for dimensionality reduction and a 3×3 convolution for feature enhancement. To guide the embedding process, the TSEAM is applied globally across the encoder to highlight semantically important regions, while the SIMAM is applied after each DenseBlock to refine local spatial features. The combination of TSEAM and SIMAM enhances embedding quality by jointly optimizing global semantics and local detail preservation. After the three-stage fusion, the encoder applies spatial downsampling to compress the feature maps, reducing their dimensions while retaining critical information. The final compressed features generate the watermark difference map (wm_diff), which is added to the input image to complete the watermark embedding process. Watermark embedding is achieved using the following formula:
$\boldsymbol{U_{\mathit{embed}}=U+\alpha\cdot\mathit{M}\cdot\mathit{wm}_{\mathit{diff}}}$ (1)
@@
In Formula (1), $U$ represents the original high-frequency sub-band, $\alpha$ controls the watermark embedding strength, M is the mask generated by the attention mechanism to weight the watermark embedding area, and $\text{wm}_{\text{diff}}$ is the generated watermark difference map. Through this process, the watermark information is embedded into the original high-frequency sub-band, forming a watermarked high-frequency sub-band $U_{\text{embed}}$. The post-processing module uses the inverse dual-tree complex wavelet transform to reconstruct the watermarked U-channel, merges it with the Y and V channels to form watermarked YUV images, and converts them to RGB images $I_{w}$, completing the embedding process.
III-B4 Noise Pool Layer (NPL)
@@ III-B5 Separable Decoder
III-B6 Structural Consistency Graph Neural Network (SC-GNN))
@@ IV-A Implementation Details
Datasets. This research is primarily based on the high-quality CelebA-HQ[43, 44]dataset, which contains 30,000 facial images and 6,217 unique identities. According to the official partitioning, it is divided into training, validation, and testing sets, where the training set contains 24,183 images, and the validation and test sets contain 2,993 and 2,824 images, respectively. In addition, to evaluate the generalization performance on out-of-distribution data, we further test the model on the LFW[45] dataset, which contains 13,233 unconstrained facial images from 5,749 identities. To verify the adaptability and robustness of the algorithm across different resolutions and simulate the diverse requirements of practical application scenarios, the experiment uses two types of images with resolutions of 128×128 and 256×256 for testing.
Fig. 6: More malicious distortions. From top to bottom are the origin image $I_{o}$, the watermarked image $I_{w}$, the noised image $I_{n}$, the residual signal $|I_{o}-I_{w}|$, and $|I_{n}-I_{w}|$. The columns indicate different types of transformations, including SimSwap, GANimation, and various StarGAN attribute modifications (Male, BlondHair, BlackHair, BrownHair, Young). All images are resized to $256\times 256$ for consistency.
Manipulation Pools. In order to comprehensively evaluate the robustness and adaptability of the algorithm under different distortion conditions, this study constructs two types of operation pools: the Common Pool and Malicious Pool. The Common Pool includes a variety of common image distortion operations, such as {Identity, JpegTest, Resize, GaussianBlur, MedianBlur, Dropout, SaltPepper, GaussianNoise} . These operations are used in the training and testing stages to simulate the general distortions that may occur in real-world scenarios, while ensuring that the facial content of the image remains unchanged. The Malicious Pool includes two common types of deepfake technologies: face swap and face replay. Specifically, SimSwap[2] is used for face swap, while GANimation[4] and StarGAN[3] are used for face replay. In the SimSwap task, the target face is randomly selected from the CelebA verification set. ”In the tasks of GANimation and StarGAN, the target expressions and attributes are randomly selected. The attributes are chosen from the set Male, Young, BlackHair, BlondHair, BrownHair. All forged images are generated using the publicly released pre-trained models to ensure the high credibility and diversity of deepfake images.
TABLE I: objective visual quality evaluation of watermarked images on celeba-hq: model, image size, message length, psnr, and ssim
@@
Parameters. This research is based on PyTorch[46] and runs on an NVIDIA RTX 4090 GPU. The Adam optimizer[47]is initially set with a learning rate of 0.0005, which is later dynamically adjusted during training. The total loss function is composed of multiple loss terms, which are combined by weighting to balance the influence of each loss on model optimization. As shown in Formula (7), during the training process, the weight parameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to 0.01, 1, 10, 10, and 10, respectively. These values are based on the preliminary experimental results and the empirical setting of the contribution of each task target to the overall optimization, aiming at achieving a good balance and stable convergence between embedding quality, detection ability and structural consistency.
TABLE II: Robustness evaluation of watermark recovery on CelebA-HQ under common and deepfake distortions at 128$\times$128 and 256$\times$256 resolutions, measured by bit error rate (BER).
@@
Baselines. To verify the performance of the proposed method, we selected several representative watermarking and proactive watermarks frameworks as baselines. Although this research belongs to the field of proactive watermarks, due to the limited existing methods and the scarcity of open source codes, we also refer to the robust watermarking method to evaluate its stability and extraction performance. In the watermarking domain, three robust watermarking methods were selected: MBRS [27], CIN [28], and PIMOG [37] (suitable for 128×128 images). In the proactive watermarks domain, we include FaceSigns[16] (for 256×256 images) and SepMark[18] (which supports both 128×128 and 256×256 resolutions). All the other baseline methods use publicly available pre-trained model weights to ensure fairness.
@@ IV-B Visual Quality
To evaluate the visual quality of the image after watermark embedding, this paper adopts two indicators: average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM)[48], which respectively quantify the changes in image signal distortion and structural fidelity, fully reflecting the impact of embedding on the image. TABLEI quantitatively verifies that our method has the best visual quality in all baselines and shows better invisibility. Experimental results show that our method achieves the highest PSNR and SSIM values at resolutions of 128×128 and 256×256. This improvement is primarily attributed to SC-GNN and attention-enhanced embedding. In addition, Fig.5 and 6 show the subjective visual contrast results at different resolutions, respectively, indicating that it is almost difficult to distinguish the watermarked image from the original image under common distortion and deepfake attacks. Combining the objective evaluation results in TABLEI with the subjective visual comparisons in Fig. 5 and 6, it is evident that the proposed method achieves superior imperceptibility in both quantitative metrics and human visual perception.
@@ IV-C Robustness Test
TABLE III: Robustness test for the image $I_{n}$ under common and malicious distortions on cross-dataset LFW
128 $\times$ 128 256 $\times$ 256
Distortion PSNR = 46.4265 PSNR = 43.7051
Tracer Detector Tracer Detector
Identity 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.0027% 0.0013% 0.0031% 0.0011%
Resize 0.0000% 0.0000% 0.0000% 0.0000%
GaussianBlur 0.0046% 0.0038% 0.0052% 0.0064%
MedianBlur 0.0023% 0.0011% 0.0011% 0.0008%
Dropout 0.0065% 0.0096% 0.0023% 0.0032%
SaltPepper 0.0116% 0.0223% 0.0043% 0.0066%
GaussianNoise 0.0319% 0.0428% 0.0087% 0.0072%
SimSwap 0.0352% 47.5363% 0.0057% 46.5418%
GANimation 0.0009% 47.5368% 0.0000% 46.5432%
StarGAN (Male) 0.0047% 47.5313% 0.0007% 46.5461%
StarGAN (Young) 0.0038% 47.5356% 0.0012% 46.5431%
StarGAN (BlackHair) 0.0096% 47.5361% 0.0008% 46.5432%
StarGAN (BlondHair) 0.0029% 47.5368% 0.0014% 46.5434%
StarGAN (BrownHair) 0.0072% 47.5363% 0.0007% 46.5468%
Average Common 0.0075% 0.0101% 0.0031% 0.0032%
$\uparrow$0.0028$\uparrow$0.0036$\uparrow$0.0014$\uparrow$0.0009%
Average Malicious 0.0092% 47.5342% 0.0015% 46.5439%
$\uparrow$0.0027—- $\downarrow$0.0007—-
TABLE IV: ablation results for psnr, ssim, and tracer’s bit error rate (%), where ‘×’ denotes non-usage and ‘$\sqrt{}$’ denotes usage.
@@
TABLE V: ablation results for psnr, ssim, and detector’s bit error rate (%), where ‘×’ denotes non-usage and ‘$\sqrt{}$’ denotes usage.
@@
TABLE II quantitatively compares the BER performance of different methods under common distortion and deepfake forgery, covering 128×128 and 256×256 resolutions. Experimental results show that our method achieves the lowest BER in all test scenarios, showing excellent robustness. Under ordinary distortion, the BER of Tracer and Detector are both close to 0%, which ensures that the watermark can be recovered with high precision under all kinds of interference. However, the BER of the contrast method is obviously higher under various common distortion conditions, which indicates that its robustness is weak. Under the deepfake forgery, the BER of Tracer is still close to 0%, while the BER of semi-robust Detector is maintained close to 50%, which effectively distinguishes the true and forged contents. Although our proposed semi-robust Detector exhibits a slightly lower BER than SepMark under deepfake attacks, the primary objective of the Detector is to distinguish between authentic and forged content. As long as a significant BER gap exists between common distortions and deepfake forgeries, effective detection can be achieved. Our method yields a substantially lower BER under common distortions, thereby enlarging the separation margin between real and fake samples and significantly enhancing discriminability. However, the existing methods have large fluctuations in BER and lack of detection reliability.
To evaluate the cross-dataset generalization ability, TABLE III reports the BER performance of our method on the LFW dataset under both common and malicious distortions at two resolutions. The values in bold represent the absolute difference in BER between LFW and the original dataset (CelebA-HQ), with arrows indicating the direction of change—$\uparrow$ for increased BER and $\downarrow$ for decreased BER on LFW. Despite the domain shift, our method maintains low BER in tracing and preserves the 50% boundary in detection across most distortion types. This suggests that our framework is not only robust under in-domain perturbations but also generalizes well to unseen data from different distributions, demonstrating strong transferability and practical applicability.
This robustness is mainly due to the frequency domain embedding strategy. Compared with the spatial domain method, the frequency domain watermark is more stable under common distortion, ensuring the integrity of the watermark information. At the same time, the design of joint sub-band enhances the discrimination ability of semi-robust Detector under deepfake forgery, so that it can stably distinguish real and forged content.
@@ IV-D Ablation Study
In order to evaluate the contribution of Joint Sub-bands, Attention Module, Frequency Domain Embedding, and GNN in this method, we conducted ablation experiments on both Tracer and Detector. The experiments were conducted on images with a resolution of 128×128. (a) Adjust the sub-bands usage of Tracer and Detector: In Tracer, watermark embedding was reduced from 1,3,4,6 sub-bands to 1,3, while in Detector, the detection was expanded from 1,3 sub-bands to 1,3,4,6. This allows us to evaluate the influence of joint sub-band information redundancy on watermark recovery and deepfake detection. (b) Remove the attention mechanism in the process of watermark embedding to analyze its influence on visual quality. (c) When the DT-CWT frequency domain embedding is removed, the watermark is directly embedded in the U channel. (d) Removing SC-GNN to evaluate its impact on the visual quality of the watermarked image. TABLE V and V present the results of the ablation study.